OpenStudy (anonymous):

Two balls of equal masses are thrown upwards along the same vertical direction at an interval of 2 s,with the same initial velocity of 39.2 m/s. The two balls will collide at a height of (a) 39.2 m (b) 73.5 m (c) 78.4 m (d) 117.6 m

OpenStudy (irishboy123):

the masses don't matter! more importantly, do you know the basic equations of motion?? here you can use \( x_1(t) = ut - \frac{1}{2} g t^2\) for the second ball, you'd need to subby in the time delay so \( x_2(t) = u(t-2) - \frac{1}{2} g (t-2)^2\) then solve out \(x_1 = x_2\). https://en.wikipedia.org/wiki/Equations_of_motion

OpenStudy (anonymous):

Thank you for answering, yes i know eqns of motion but did not know how to apply them here. The answer given is (b)73.5 , but still not able to solve it. Can you please explain in a detailed manner.